Joe has a collection of nickels and dimes that is worth $6.05. If the number of dimes was doubled and the number of nickels was decreased by 10, the value of the coins would be $9.85. How many dimes does he have?

Answer by stanbon(75887)

Joe has a collection of nickels and dimes that is worth $6.05. If the number of dimes was doubled and the number of nickels was decreased by 10, the value of the coins would be $9.85. How many dimes does he have?
-------------------------
Value Equation: 5n + 10d = 605 cents
Value Equation: 10(2d) + 5(n-10) = 985 cents
----------------------------------
Modified Equations:
20d + 5n = 1035
10d + 5n = 605
-------------------
10d = 430
d = 43 (# of dimes)
---
Substitute into 5n + 10d = 605 to solve for "n":
5n + 10*43 = 605
5n = 175
n = 35 (# of nickels)
------------------------------
Cheers,
Stan H.